What is the strength (% w/w) of a cortisone ointment prepared by mixing 50 grams of 8% ointment, 100 grams of 12% cort...

  1. What is the strength (% w/w) of a cortisone ointment prepared by mixing 50 grams of 8% ointment, 100 grams of 12% cortisone ointment, and 20 grams of ointment base?

    5.00%

    9.33%

    8.57%

    9.41%

    6.67%
Answer #1

Total mass = 50+100+20 = 170 grams.

Mass of cortisone oinment = 12×100/100 = 12 grams

Mass ointment = 8×50/100 = 4 gram

Total cortisone ointment = 12+4 = 16 gram

%w/w of cortisone = mass of cortisone ointment×100/total mass

= 16×100/170

= 9.41 %